OpenStudy (anonymous):

The amplitude and period of y = cos3xcos7x + sin3xsin7x are ?

OpenStudy (callisto):

y = cos3xcos7x + sin3xsin7x = cos (3x-7x) = cos (-4x) = cos4x For amplitude, I think it should be the same as the ordinary case?! For period, i think it would be reduced by 1/4, that is pi/2 ?! I'm not sure...

OpenStudy (anonymous):

Amp = 1 and Period = pi/2 so you got the period correct but i dont get the Amp though ?

OpenStudy (callisto):

there is no change in amplitude as there is no enlargement along y-axis.. If you see it is like y = kf(x) there is some change in the amplitude. But for this case, it's not